OpenStudy (turingtest):

Let\[x_1,x_2,...\]be a succession of real, non-negative numbers such that\[x_{n+1}\le x_n+\frac{1}{n^2}\]for all\[n\ge1\]Show that\[\lim_{n \rightarrow \infty}x_n\]exists.

OpenStudy (turingtest):

Time for breakfast first...

OpenStudy (zarkon):

*bookmark (I'll be on later)

OpenStudy (zarkon):

ha...trivial...I'll let you think about it :)

OpenStudy (turingtest):

That's only fair, y'all basically solved the other two for me. I need to learn to think about these kinds of problems.

OpenStudy (asnaseer):

try writing out the terms TuringTest - you should see how to solve it then...

OpenStudy (asnaseer):

use the fact that the MAXIMUM \(x_{n+1}=x_n+\frac{1}{n^2}\)

OpenStudy (turingtest):

Okay, this kinda thing is really new to me, so I'm gonna be a bit slow getting it, but thanks for the tip!

OpenStudy (asnaseer):

what would the first term be?

OpenStudy (turingtest):

It doesn't say a specific number, it's just a succession of positive real numbers. So you just mean\[x_2\le x_1+1\]eh?

OpenStudy (asnaseer):

yes - first term is \(x_1\) 2nd term MAX is \(x_1+\frac{1}{1^2}\) now, what would the 3rd term MAX be?

OpenStudy (turingtest):

obviously\[x_2+\frac{1}{4}\]

OpenStudy (asnaseer):

don't add up the values - leave them as 1/n^2

OpenStudy (asnaseer):

and write the MAX 3rd term in terms of \(x_1\)

OpenStudy (asnaseer):

you should spot a pattern emerging...

OpenStudy (turingtest):

ah... x_3 max is\[x_2+\frac{1}{2^2}=x_1+1+\frac{1}{2^2}\]x_n max is\[x_{n-1}+1+\frac{1}{2^2}+\dots+\frac{1}{(n-1)^2}\]

OpenStudy (asnaseer):

1st term should be \(x_1\)

OpenStudy (turingtest):

yeah, just didn't write it. \[x_1\]x_2 max is\[x_2=x_1+1\]x_3 max is\[x_2+\frac{1}{2^2}=x_1+1+\frac{1}{2^2}\]x_n max is\[x_{n-1}+1+\frac{1}{2^2}+\dots+\frac{1}{(n-1)^2}\]

OpenStudy (asnaseer):

you should end up with something like:\[\text{MAX }x_{n+1}=x_1+\sum_1^n\frac{1}{n^2}\]

OpenStudy (asnaseer):

the sum on the right has a well known value

OpenStudy (turingtest):

typo again... \[x_1\]x_2 max is\[x_2=x_1+1\]x_3 max is\[x_2+\frac{1}{2^2}=x_1+1+\frac{1}{2^2}\]x_n max is\[x_1+1+\frac{1}{2^2}+\dots+\frac{1}{(n-1)^2}\]so that leads to your formula clearly

OpenStudy (asnaseer):

BTW: I meant the 1st in the last line of your reply should be \(x_1\)

OpenStudy (asnaseer):

ok - so do you recognise the sum?

OpenStudy (turingtest):

yeah, caught that^^^ server is so slow...

OpenStudy (asnaseer):

remember we want to know what happens as n tends to \(\inf\)

OpenStudy (asnaseer):

actaully we just need to show that a limit exists - you don't need the value of it

OpenStudy (jamesj):

Ok, so the x_n are bounded by \( x_1 + π^2/6 \). Now why does that mean they have a limit?

OpenStudy (turingtest):

Well because it's finite of course

OpenStudy (asnaseer):

the sum is well known and has a value of \(\frac{\pi^2}{6}\) as JamesJ stated

OpenStudy (turingtest):

That I couldn't recall, though I knew it was finite

OpenStudy (asnaseer):

so we now know that:\[x_{n+1}\le x_1+\frac{\pi^2}{6}\]

OpenStudy (turingtest):

isn't that enough to say that the limit is finite because x_1 and the sum are finite? regardless of n?

OpenStudy (turingtest):

and that the sum exists of course as James pointed out.

OpenStudy (asnaseer):

maybe this is not a /strong/ enough proof. I thought this was enough but there must be more to it. maybe @JamesJ can help us understand?

OpenStudy (jamesj):

The thing is you can have a bounded sequence that doesn't have a limit. For instance a_n = (-1)^n What you can say is that a bounded sequence has a convergent subsequence. You might want to try and show the sequence converges to the limit of a subsequence. But you may not want to go down that road.

OpenStudy (turingtest):

Oh man, I don't even know what a subsequence is. Is that the sequence of partial sums?

OpenStudy (asnaseer):

do we need to use L'Hôpital's rule?

OpenStudy (jamesj):

there are no sums here, except in the bound calculation we just did.. You want to show that (xn) converges. Not the sum of (xn) or anything else.

OpenStudy (asnaseer):

do we just need to show that the ratio between successive terms gets smaller, and therefore it converges?

OpenStudy (jamesj):

(And L'Hopital has absolutely nothing to do with it.)

OpenStudy (asnaseer):

sorry - just grasping at straws here :-(

OpenStudy (jamesj):

A strategy would be to show that (xn) is ultimately a Cauchy sequence. Then (xn) converges.

OpenStudy (jamesj):

These terms--subseqnece, Cauchy--are all from Real Analysis. Having some knowledge of these things and standard theorems will help.

OpenStudy (jamesj):

*subsequence

OpenStudy (asnaseer):

\[x_{n+1}-x_n\le\frac{1}{n^2}\]and we know 1/n^2 converges - is that sufficient?

OpenStudy (turingtest):

the Cauchy sequence approach is what asnaseer said isn't it? Showing that the elements get arbitrarily closer to each other as n increases

OpenStudy (jamesj):

If you absolute value then yes, that would show (xn) is Cauchy and then you could conclude (xn) converges. But unfortunately, you don't have \[ | x_{n+1} - x_n | < 1/n^2 \]

OpenStudy (turingtest):

:(

OpenStudy (asnaseer):

JamesJ - does it have to be strictly LESS THAN? LESS THAN OR EQUAL is not enough?

OpenStudy (turingtest):

right because it can be large negative...

OpenStudy (jamesj):

less than/less than or equal; doesn't matter. What's important is the absolute value. Anyway, I think the Cauchy argument is the way to go. Suppose the (xn) is not Cauchy. Write that down in \( \epsilon-N \) terms and derive a contradiction.

OpenStudy (asnaseer):

ok - thanks - time to go learn about Cauchy...

OpenStudy (turingtest):

Yikes, I haven't done that sort of thing yet. Yes, study time!

OpenStudy (asnaseer):

which is good - we are increasing our knowledge! :-)

OpenStudy (turingtest):

That's why I'm here :-)

OpenStudy (asnaseer):

ok, I've learned a little about Cauchy Sequences and Real Analysis (quite interesting) and have formulated this so far:\(\{x_n\}\) is Cauchy if for any \(\epsilon>0\) there is a number N such that \(|x_n-x_m|<\epsilon\) for n,m>N. So we can write:\[\begin{align} |x_n-x_m|&=|(x_n-x_{n-1})+(x_{n-1}-x_{n-2})+...+(x_{m+1}-x_m)|\\ &\le|x_n-x_{n-1}|+|x_{n-1}-x_{n-2}|+...+|x_{m+1}-x_m|\\ &\le\frac{1}{(n-1)^2}+\frac{1}{(n-2)^2}+...+\frac{1}{m^2}\\ &\le\psi^1(m)-\psi^1(n+1) \end{align}\] I'm not sure how to proceed from here. Am I even on the right track here @JamesJ?

OpenStudy (jamesj):

As we commented above, you don't know that \[ |x_n - x_{n-1} | \leq \frac{1}{(n-1)^2} \] If you did, this problem would be easy. But alas, we don't.

OpenStudy (turingtest):

Why did Zarkon call this problem 'trivial' if it involves so much I wonder...

OpenStudy (asnaseer):

but we are given:\[x_{n+1}\le x_n+\frac{1}{n^2}\]therefore:\[x_{n+1}-x_n\le \frac{1}{n^2}\]and since the right hand side is always positive aren't we allowed to just add the |'s around the left hand side?

OpenStudy (asnaseer):

@Zarkon - have we missed one of your BIG FOOTPRINTS somewhere?

OpenStudy (zarkon):

there is more than one way to do a problem

OpenStudy (jamesj):

It's actually not hard if you know how to write down the converse of the Cauchy condition. But perhaps he has something else in mind.

OpenStudy (asnaseer):

@Zarkon - I think your reply should have been: "to do a problem, more than one way, there is" in the style of Yoda :-)

OpenStudy (zarkon):

I'll keep that in mind ;)

OpenStudy (asnaseer):

can you give us a clue @Zarkon?

OpenStudy (asnaseer):

@JamesJ - why can't we write the abs value as indicated?

OpenStudy (jamesj):

Suppose x_4 = 1. Then \[ x_5 \leq x_4 + 1/4^2 = 17/16 \] so \( x_5 \) could be 1/2 say, in which case it is not true that \[ | x_5 - x_4 | \leq 1/4^2 \]

OpenStudy (asnaseer):

so for recursive definitions of inequalities - we can't use the same rules as for normal inequalities I guess

OpenStudy (zarkon):

consider the sequence defined by \[y_1=x_1\] and \[y_{n+1}=x_{n+1}-\sum_{i=1}^{n}\frac{1}{i^2}\]

OpenStudy (jamesj):

Nice.

OpenStudy (asnaseer):

\(y_{n+1}\) is always equal to \(x_1\)?

OpenStudy (zarkon):

no

OpenStudy (asnaseer):

sorry always <= x_1

OpenStudy (zarkon):

\[y_{n+1}=x_{n+1}-\sum_{i=1}^{n}\frac{1}{i^2}\le x_{n}+\frac{1}{n^2}-\sum_{i=1}^{n}\frac{1}{i^2} \] \[=x_{n}-\sum_{i=1}^{n-1}\frac{1}{i^2}=y_n\]

OpenStudy (jamesj):

i.e., \( y_{n+1} \leq y_n \). Now ...

OpenStudy (asnaseer):

so the {\(y_n\)} sequence converges

OpenStudy (zarkon):

yes...because it is bonded below and decreasing.

OpenStudy (zarkon):

*bounded

OpenStudy (asnaseer):

and there must be a method to show that the {\(x_n\)} sequence also converges because of this - more reading?

OpenStudy (zarkon):

let \[y_n\to y\] \[\lim_{n\to \infty}x_n=\lim_{n\to \infty}y_{n}+\sum_{i=1}^{n-1}\frac{1}{i^2}=y+\frac{\pi^2}{6}\]

OpenStudy (jamesj):

I'll be honest and say I prefer this method to mine because it's direct; vs. mine which would be an indirect proof (and theoretically, a little more difficult).

OpenStudy (turingtest):

Well I'm lagging pretty far behind the 3 of you, guess I don't have much to contribute. I did at least follow it though up until that last step. What is y?

OpenStudy (asnaseer):

thats a nice proof - thanks @Zarkon. and thanks @JamesJ for introducing me to the world of Real Analysis.

OpenStudy (zarkon):

the limit of \(y_n\) we do now know its value

OpenStudy (turingtest):

ah, ok, then I guess I get it. I'll have to study it though. Thanks guys.

OpenStudy (jamesj):

The result that was just used is one of the most important basic results in analysis and worth knowing: Let (x_n) be sequence of real numbers which is - increasing and bounded above; or - decreasing and bounded below then (x_n) has a limit.

OpenStudy (asnaseer):

thx @JamesJ